OpenStudy (anonymous):

True or False? The point-slope form of a line is y + y1 = m(x + x1), where m is the slope and (x1, y1) is a point on the line. It's false right?

OpenStudy (anonymous):

Why would you say it's false?

OpenStudy (anonymous):

I'm thinking about y2 - y1 x2 - x1

OpenStudy (anonymous):

That's how you get slop

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

m = y2-y1/x2-x1

OpenStudy (anonymous):

oh so that statement about the question is true?

OpenStudy (anonymous):

If i gave you the point slope form

OpenStudy (anonymous):

and told you to use (4,2),(5,3) what would you do with the point slope form?

OpenStudy (anonymous):

I want you to understand this first, before you pick your answer.

OpenStudy (anonymous):

What are x1 and y1

OpenStudy (anonymous):

how would you incorporate them in the point slope form?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

Chinagurl try what I gave you.

OpenStudy (anonymous):

y1 = 2 y2 =3 x1 = 4 x2 = 5

OpenStudy (anonymous):

ok so now what do you do with that information?

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

plug in the point slope form

OpenStudy (anonymous):

How though

OpenStudy (anonymous):

but what happened to x2 & y2?

OpenStudy (anonymous):

Show me what you would with those numbers in that form

OpenStudy (anonymous):

ok so you see that m = slope right?

OpenStudy (anonymous):

m= rise/run = y2-y1/x2-x1

OpenStudy (anonymous):

y - 2 = (x - 4)

OpenStudy (anonymous):

where is your slope?

OpenStudy (anonymous):

m = 3 - 2/ 5 - 4

OpenStudy (anonymous):

Yes and you got 1 right?

OpenStudy (anonymous):

so slope m = 1 and you got y-2=(x-4)

OpenStudy (anonymous):

I'm not sure if you've done anything passed that but you'd want to put it in y=mx+b form.

OpenStudy (anonymous):

But good job!

OpenStudy (anonymous):

yes! I got one. meaning the question is true

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

thanks batman for coming to my rescue :)